OpenStudy (anonymous):

is my answer right?? Simplify: x^2 - 15x + 56 over x - 7 answer: x^2 - 23

OpenStudy (anonymous):

r u dividing it ?

OpenStudy (anonymous):

or multiplied

OpenStudy (anonymous):

subtracting

OpenStudy (anonymous):

x^2-14x+49

OpenStudy (anonymous):

x^2-15x+56/x-7 =(x-8)(x-7)/(x-7) answer= x-8

OpenStudy (anonymous):

i got x + 8

OpenStudy (anonymous):

i am getting x-8. Conform the answer from someone else too.